# Step-back Profiling: Distilling User History for Personalized Scientific Writing

### v1

@@ Abstract
Large language models (LLMs) excel at a variety of natural language processing tasks, yet they struggle to generate personalized content for individuals, particularly in real-world scenarios like scientific writing. Addressing this challenge, we introduce Step-back Profiling to personalize LLMs by distilling user history into concise profiles, including essential traits and preferences of users. Regarding our experiments, we construct a Personalized Scientific Writing (PSW) dataset to study multi-user personalization. PSW requires the models to write scientific papers given specialized author groups with diverse academic backgrounds. As for the results, we demonstrate the effectiveness of capturing user characteristics via Step-back Profiling for collaborative writing. Moreover, our approach outperforms the baselines by up to 3.6 points on the general personalization benchmark (LaMP), including 7 personalization tasks. Our extensive ablation studies validate the contributions of different components in our method and provide insights into our task definition. Our dataset and code are available at https://github.com/gersteinlab/step-back-profiling.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have made significant progress in natural language understanding and generation (Wei et al., 2022a; Zhang et al., 2023b; OpenAI, 2023; Qin et al., 2023). Concurrently, integrating LLMs with personalization paradigms has paved the way for a vast frontier in improving user-centric services and applications (Salemi et al., 2023; Chen et al., 2023; Zhiyuli et al., 2023), as they provide a deeper understanding of users’ accurate demands and interests than abstract vector-based information representations. By learning to characterize and emulate user-specific language patterns, personalized LLMs can enable more engaging and valuable interactions in domains such as dialogue (Wang et al., 2019; Zhang et al., 2019b; Character.AI, 2022), recommendation (Zhiyuli et al., 2023; Wang et al., 2023a), role-playing (Shao et al., 2023; Jiang et al., 2023) and content creation (Cao et al., 2023; Wei et al., 2022b).
Figure 1: Overview of the Step-back Profiling.
@@
Figure 2: Step-back Profiling performance on the LaMP benchmark. Details of experimental setup can be found in Section 4.2.
@@ 2.2 Procedure
Each user’s history is condensed into a short “gist” representation using an abstraction function $\operatorname{Gist}(\cdot)$: $\mathbb{P}_{u_{i}}=\operatorname{Gist}(\mathbb{H}_{i})$. The “gist” captures the user’s high-level traits and interests.
@@
Individual user profiles $\mathbb{P}_{u_{1}}$, $\mathbb{P}_{u_{2}}$, $\cdots$, $\mathbb{P}_{u_{n}}$ are concatenated to form a unified representation $\mathbb{P}_{\mathbb{U}}=[\mathbb{P}_{u_{1}};\mathbb{P}_{u_{2}};\cdots;\mathbb{P}_{u_{n}}]$, where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is a permutation-sensitive function combining the user profiles.
@@
Relevant snippets from user histories $\mathbb{H}_{\mathbb{U}}$ may be retrieved for input $x$ using a retrieval function $\operatorname{Retrieve}(\cdot)$. We have $\mathbb{R}_{i,k}=\operatorname{Retrieve}(x,\mathbb{H}_{i},k)$, where $\mathbb{R}_{i,k}$ is a set of top-$k$ retrieved input-output snippets from user $u_{i}$’s history $\mathbb{H}_{i}$. The top-$k$ retrieved snippets $\mathbb{R}_{k}=\{\mathbb{R}_{i,k}\}_{i=1}^{N}$ can be concatenated with $x$ to form an augmented input $\hat{x}=[x;\mathbb{R}_{1,k};\mathbb{R}_{2,k};\cdots;\mathbb{R}_{n,k}]$.
@@
The personalized language model generates an output $y=\operatorname{Generate}(\hat{x},\mathbb{P}_{\mathbb{U}})$ by conditioning on the augmented input $\hat{x}$ (if retrieval is used) or the original input $x$, along with the concatenated user profile $\mathbb{P}_{\mathbb{U}}$. The generated output $y$ aligns with the user preferences captured by the Step-back Profiling while following the input $x$.
@@ 3.3 GPT-based Evaluation
LLM-based evaluators, such as G-Eval, have shown high consistency with human evaluators Liu et al. (2023); Chang et al. (2024), particularly in personalized text generation Wang et al. (2023b). Therefore, we utilize GPT-4-turbo with chain-of-thought prompting as a judge to evaluate the generated outputs on the PSW benchmark in multiple dimensions (Zhang et al., 2019a), including consistency, fluency, relevance, and novelty. An example of our evaluation (G-Eval) prompt can be found in Appendix C.
@@ 4.2 Main Result
    Single-Author: Personalizes with single user’s profile $P_{u_{i}}$ and retrieved snippets $R_{i}$: $y=\operatorname{Generate}(\hat{x},P_{u_{i}})$, where $\hat{x}=[x;\mathbb{R}_{i}]$ and $\mathbb{R}_{i}=\operatorname{Retrieve}(x,\mathbb{H}_{i},10)$.
@@
    Multi-Author: Personalizes with multiple users’ profiles $\mathbb{P}_{\mathbb{U}}$ and retrieved snippets $\mathbb{R}$: $y=\operatorname{Generate}(\hat{x},\mathbb{P}_{\mathbb{U}})$, where $\hat{x}=[x;\mathbb{R}_{1};\cdots;\mathbb{R}_{n}]$, $\mathbb{R}_{i}=\operatorname{Retrieve}(x,\mathbb{H}_{i},10)$ for each user $u_{i}$.
@@ 5 Conclusion
Scalability and Efficiency Although the Step-back Profiling method improves memory management, the approach still has scalability concerns, particularly with very large user histories or an increasing number of collaborators. Efficiently managing and retrieving relevant user data from extensive histories without compromising performance remains a challenge.
@@ Appendix D Prompts for LaMP Tasks
D.1 Personalized Citation Identification (LaMP-1)
@@
D.2 Personalized News Categorization (LaMP-2)
@@
D.3 Personalized Product Rating (LaMP-3)
@@
D.4 Personalized News Headline Generation (LaMP-4)
@@
D.5 Personalized Scholarly Title Generation (LaMP-5)
@@
D.6 Personalized Email Subject Generation (LaMP-6)
@@
D.7 Personalized Tweet Paraphrasing (LaMP-7)

### v2

@@ Abstract
Large language models (LLM) excel at a variety of natural language processing tasks, yet they struggle to generate personalized content for individuals, particularly in real-world scenarios like scientific writing. Addressing this challenge, we introduce Step-back Profiling to personalize LLMs by distilling user history into concise profiles, including essential traits and preferences of users. To conduct the experiments, we construct a Personalized Scientific Writing (PSW) dataset to study multi-user personalization. PSW requires the models to write scientific papers given specialized author groups with diverse academic backgrounds. As for the results, we demonstrate the effectiveness of capturing user characteristics via Step-back Profiling for collaborative writing. Moreover, our approach outperforms the baselines by up to 3.6 points on the general personalization benchmark (LaMP), including 7 personalization LLM tasks. Our ablation studies validate the contributions of different components in our method and provide insights into our task definition. Our dataset and code are available at https://github.com/gersteinlab/step-back-profiling.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have made significant progress in natural language understanding and generation (Wei et al., 2022a; Zhang et al., 2023b; OpenAI, 2023; Qin et al., 2023). Concurrently, integrating LLMs with personalization paradigms has paved the way for a vast frontier in improving user-centric services and applications (Salemi et al., 2023; Chen et al., 2023; Zhiyuli et al., 2023), as they provide a deeper understanding of users’ accurate demands and interests than abstract vector-based information representations. By learning to characterize and emulate user-specific language patterns, personalized LLMs can enable more engaging and valuable interactions in domains such as dialogue (Wang et al., 2019; Zhang et al., 2019b; Character.AI, 2022), recommendation (Zhiyuli et al., 2023; Wang et al., 2023a), role-playing (Shao et al., 2023; Jiang et al., 2023; Tang et al., 2024) and content creation (Wei et al., 2022b; Cao et al., 2023).
Figure 1: Overview of the Step-back Profiling. Step-back Profiling creates compact profile representations for each user and augments LLM generation with personalized information.
@@
Figure 2: Step-back Profiling performance on the LaMP benchmark. We observe dominant performance of Step-back Profiling on most LaMP tasks. Details of experimental setup can be found in Section 4.2.
In the realm of scientific publishing, we observe two distinct paradigms: papers authored by individuals and those resulting from team science collaborations. The latter category, exemplified by consortium science projects, can involve hundreds of authors working together (Wuchty et al., 2007). This shift towards large-scale collaborative research necessitates a deeper understanding of how to effectively model and generate content that reflects the collective expertise and writing styles of multiple authors. Our work aims to address this gap by developing methods to mimic and enhance the collaborative writing process in team science scenarios.
@@
By focusing on multi-author personalization, our work not only addresses a critical gap in current LLM research but also aligns with the evolving landscape of scientific collaboration. The PSW dataset and our Step-back Profiling approach provide a foundation for understanding and enhancing the complex dynamics of team-based scientific writing, potentially leading to more efficient and effective collaborative research practices.
@@ 2.2 Procedure
Each user’s history is condensed into a short “gist” representation using an abstraction function $\operatorname{Gist}(\cdot)$: $\mathbb{P}_{u_{i}}=\operatorname{Gist}(\mathbb{H}_{i})$. The “gist” captures the user’s high-level traits and interests. In Step-back Profiling, the $\operatorname{Gist}(\cdot)$ is also achieved using a language model. Details of the prompts can be found in Appendix D and E.
@@
Individual user profiles $\mathbb{P}_{u_{1}}$, $\mathbb{P}_{u_{2}}$, $\cdots$, $\mathbb{P}_{u_{n}}$ are concatenated to form a unified representation $\mathbb{P}_{\mathbb{U}}=[\mathbb{P}_{u_{1}};\mathbb{P}_{u_{2}};\cdots;\mathbb{P}_{u_{n}}]$, where [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is an order-sensitive function combining the user profiles.
@@
Relevant snippets from user histories $\mathbb{H}_{\mathbb{U}}$ may be retrieved for input $x$ using a retrieval function $\operatorname{Retrieve}(\cdot)$. We have $\mathbb{R}_{i,k}=\operatorname{Retrieve}(x,\mathbb{H}_{i},k)$, where $\mathbb{R}_{i,k}$ is a set of top-$k$ retrieved input-output snippets from user $u_{i}$’s history $\mathbb{H}_{i}$. The top-$k$ retrieved snippets $\mathbb{R}_{k}=\{\mathbb{R}_{i,k}\}_{i=1}^{N}$ can be concatenated with $x$ to form an augmented input $\tilde{x}=[x;\mathbb{R}_{1,k};\mathbb{R}_{2,k};\cdots;\mathbb{R}_{n,k}]$.
@@
The personalized language model generates an output $y=\operatorname{Generate}(\tilde{x},\mathbb{P}_{\mathbb{U}})$ by conditioning on the augmented input $\tilde{x}$ (if retrieval is used) or the original input $x$, along with the concatenated user profile $\mathbb{P}_{\mathbb{U}}$. The generated output $y$ aligns with the user preferences captured by the Step-back Profiling while following the input $x$.
@@ 3 The Personalized Scientific Writing (PSW) Benchmark
3.1 Motivation
The Personalized Scientific Writing (PSW) Benchmark is motivated by the evolving landscape of scientific research, where large-scale collaborations involving numerous authors across disciplines have become increasingly common (Wuchty et al., 2007; Wu et al., 2019). While existing benchmarks like LaMP (Salemi et al., 2023) focus on single-user personalization, they fail to address the complex dynamics of multi-author scenarios prevalent in modern scientific writing. PSW aims to bridge this gap by providing a framework for evaluating personalized language models in collaborative contexts, challenging them to capture and reproduce the intricate interplay of individual authors’ styles, expertise, and roles within a team (Fortunato et al., 2018).
By extending personalization to multi-user scenarios, PSW reflects real-world applications in interdisciplinary research, where integrating diverse knowledge and writing styles is crucial. The benchmark assesses how well language models can synthesize collective intelligence in writing tasks, a critical factor in enhancing the quality and impact of research outputs (Lee and Haupt, 2021). Through PSW, we aim to drive advancements in personalized language modeling techniques that can handle multiple user profiles simultaneously, potentially contributing to more efficient and impactful research practices in the era of team science and artificial intelligence-assisted discovery (King et al., 2009).
@@ 3.4 GPT-based Evaluation
LLM-based evaluators, such as G-Eval (Liu et al., 2023), have shown high consistency with human evaluators Chang et al. (2024), particularly in personalized text generation Wang et al. (2023b). Therefore, we utilize GPT-4-turbo with chain-of-thought prompting as a judge to evaluate the generated outputs on the PSW benchmark in multiple dimensions (Zhang et al., 2019a), including consistency, fluency, relevance, and novelty. An example of our evaluation (G-Eval) prompt can be found in Appendix C.
@@ 4.2 Main Result
    Single-Author: Personalizes with single user’s profile $P_{u_{i}}$ and retrieved snippets $R_{i}$: $y=\operatorname{Generate}(\tilde{x},P_{u_{i}})$, where $\tilde{x}=[x;\mathbb{R}_{i}]$ and $\mathbb{R}_{i}=\operatorname{Retrieve}(x,\mathbb{H}_{i},10)$.
@@
    Multi-Author: Personalizes with multiple users’ profiles $\mathbb{P}_{\mathbb{U}}$ and retrieved snippets $\mathbb{R}$: $y=\operatorname{Generate}(\tilde{x},\mathbb{P}_{\mathbb{U}})$, where $\tilde{x}=[x;\mathbb{R}_{1};\cdots;\mathbb{R}_{n}]$, $\mathbb{R}_{i}=\operatorname{Retrieve}(x,\mathbb{H}_{i},10)$ for each user $u_{i}$.
@@ 5 Conclusion
Scalability in Multi-Author Scenarios The current implementation of Step-back Profiling may face challenges in scenarios with a very large number of authors, as individual personalities might become diluted or “washed out” in the collective profile. This limitation could affect the effectiveness of personalization in large-scale collaborative projects. Future research should investigate the scalability of Step-back Profiling as the number of authors increases. By systematically varying the number of collaborators, we can determine how personalization effectiveness changes and at what point, if any, the benefits of individualized profiles diminish. This will help establish guidelines for applying Step-back Profiling across various collaborative scenarios.
Optimal Collaborative Writing Strategies Our current study does not explore the optimal strategies for collaborative scientific writing using Step-back Profiling. To address this limitation, we propose future research aimed at identifying the most effective approaches to team-based scientific writing. This could involve comparing various team compositions, such as groups of diverse experts versus more homogeneous teams, or analyzing different collaborative processes enabled by Step-back Profiling. The research would aim to quantify how different collaborative strategies impact the quality, innovation, and efficiency of scientific writing. Findings from this study could provide valuable insights into best practices for team formation and collaboration in scientific writing, potentially leading to more effective use of AI-assisted tools like Step-back Profiling in academic and research contexts.
@@ Appendix D Prompts for LaMP Tasks
D.1 User Profile Generation
User History
I will provide you with some examples of the user’s past interactions. Please analyze these to create a user profile. Each example consists of an input and the corresponding output.
Example 1
Input: {Input Text}
Output: {Output Text}
Example 2
Input: {Input Text}
Output: {Output Text}
Example k
Input: {Input Text}
Output: {Output Text}
Instruction
Based on the provided examples, please generate a user profile in the following format:
Keywords: [keyword1, keyword2, …]
Topics: [topic1, topic2, …]
Writing Style: [style1, style2, …]
Preferences: [preference1, preference2, …]
Table 7: Prompt template for LaMP User Profile Generation.
D.2 Personalized Citation Identification (LaMP-1)
@@
D.3 Personalized News Categorization (LaMP-2)
@@
D.4 Personalized Product Rating (LaMP-3)
@@
D.5 Personalized News Headline Generation (LaMP-4)
@@
D.6 Personalized Scholarly Title Generation (LaMP-5)
@@
D.7 Personalized Email Subject Generation (LaMP-6)
@@
D.8 Personalized Tweet Paraphrasing (LaMP-7)